OpenStudy (anonymous):

Find the points of self-intersection for the curve x(t) = t^2-2t, y(t) = t^3 - 3t^2 + 2t, t∈ℝ and then find the area of the loop.

OpenStudy (anonymous):

How does a curve exactly self intersect? >.> .

OpenStudy (anonymous):

Well, neither one will intersect with itself, but x(t) will intersect y(t). The system is a loop if you're talking about a parametric equation.

OpenStudy (anonymous):

Yeah, it's parametric.

OpenStudy (anonymous):

How would I exactly find this point? >.> .

OpenStudy (anonymous):

Whenever you have \(x(t_1) = x(t_2)\)

OpenStudy (anonymous):

\(y(t_1) = y(t_2)\)

OpenStudy (anonymous):

Umm... Okay, Just any random value of t should work right?

OpenStudy (anonymous):

It's two values of \(t\) though...

OpenStudy (anonymous):

I am thinking of setting them equal to each other?

OpenStudy (anonymous):

THis is a tricky question.

OpenStudy (anonymous):

One thing you can do is just try graphing it.

OpenStudy (anonymous):

I would normally say that you should try to isolate 't' and substitute into the other equation. However, this does not happen nicely in this case.

OpenStudy (anonymous):

Make a table of \(t\), \(x\), and \(y\) values.

OpenStudy (anonymous):

@qpHalcy0n If you know a parametrized curve self-intersects, you know it's going to be hard to de-parametrize.

OpenStudy (anonymous):

There isn't an elegant solution for this.

OpenStudy (anonymous):

I mean there isn't an elegant solution that I know of for solving self intersection among parametrized curves that I know of.

OpenStudy (anonymous):

I have to do the entire questione by hand :( .

OpenStudy (anonymous):



OpenStudy (anonymous):

Try filling this out a bit and see what happens.

OpenStudy (anonymous):

Ew... Alright. I just need a visual of this I guess. No harm in plotting then right? >.> .

OpenStudy (anonymous):

Okay, how about putting \(x(t)\) in vertex form?

OpenStudy (anonymous):

Completing the square?

OpenStudy (anonymous):

Yes, if it is in vertex form then you know the line of symmetry.

OpenStudy (anonymous):

And you can find a general equation for \(t_2\) given \(t_1\) such that \(x(t_1)=x(t_2)\)

OpenStudy (anonymous):

Well if I complete the square I get x(t) = (t-1)^2 - 1

OpenStudy (anonymous):

Okay so the vertex is at \((1, -1)\)

OpenStudy (anonymous):

Don't you mean (-1,0) ?

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

No never mind. My mistake.

OpenStudy (anonymous):

Anyway, basically if I give you \(t_1\) how would you get \(t_2\)? You reflect it over the axis of symmetry.

OpenStudy (anonymous):

Here is a visual: